OpenStudy (anonymous):

Does anyone know how to solve a question that ask's " What is the correct proportion for corresponding sides"?

OpenStudy (anonymous):

for what class?

OpenStudy (anonymous):

Geometry

OpenStudy (anonymous):

are those similar triangle problem?

OpenStudy (anonymous):

Similar Rectangles

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So if you have a rectangle with width 4 and height 2. Other similar rectangle with height 7 what is the width

OpenStudy (anonymous):

14?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Okay, the equation shows 2 triangles. A smaller one with 4cm Height & 12cm Width.. The bigger one shows 8m Height & 24m Width.. How Will i Set Up The Correct Proportion for the correspondong sides?

OpenStudy (anonymous):

are these right triangle

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Sorry, i meant to say they're rectangles

OpenStudy (anonymous):

lenght/width of one rectangle should be same as lenght/width of other

OpenStudy (anonymous):

it's Doubled

OpenStudy (anonymous):

4/12=1/3 8/24=1/3

OpenStudy (anonymous):

So 4/12 = 24/8.. That Would Be The correct Answer?

OpenStudy (anonymous):

4/12=8/24

OpenStudy (anonymous):

There is No Answer Choice For That, The Choices are 12/8=24/4 or 12/4=24/8 or 4/12=24/8 or 12/4 = 8/24.

OpenStudy (radar):

What are the choices? Is their a 1/3?